A building worth $828,000 is depreciated for tax purposes by its owner using the straight-line depreciation method. The value of the building y, after x months of use, is given by y=828,000-2200x dollars. After how many years will the value of the building be $722,400?

Answer by ewatrrr(24785)

 
Hi,
The value of the building y, after x months of use, is given by
y=828,000-2200x dollars
722400 = 828,000 - 2200x
2200x = 828,000 - 722,400
x = 105,600/2200 = 48mo 0r 4 years